3. A rubber chicken is launched straight up at speed v from ground level. Find each of the following if the launch speed is tripled (in terms of any constants and v).
a. max height
b. hang time
c. impact speed
1
Expert's answer
2021-10-04T14:01:41-0400

Given,

Initial launcher speed = v

If initial launcher speed =3v

a) Let, initially the maximum height was h, and finally it was H

Applying newton's law of motion,

"v^2= u^2-2gh"

Now, at the maximum height v=0

"\Rightarrow 0=v^2-2gh"

"h=\frac{v^2}{2gh}"

and finally "H=\frac{9v^2}{2g}"

b)

Let the time of hang is T, and time taken to reach at maximum height = t

Now, "v=u-gt"

at the top most point, final velocity of the object = 0

"\Rightarrow 0=v-gt"

"\Rightarrow t= \frac{v}{g}"

Hence, initially the time of flight of the projectile "(T)=2t = \frac{2v}{g}"

and for the final condition,

"T' = \frac{6v}{g}"


c)

The impact speed will be v if it was projected with v and it will be 3v, if it is projected with 3v.

but if "3v\ge V_{esc}" ,

then it will not come back in the gravitational field.